OpenStudy (anonymous):

STUDYING FOR MAC 1105 EXAM Solve 3x^2-8x+2=0 using the quadratic formula. Leave answers in reduced radical form.

OpenStudy (anonymous):

I have the problem worked from a test key but there are steps I don't understand

OpenStudy (anonymous):

\[ax^2+bx+c=0\to x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[ 3x^2-8x+2=0 \to x=\frac{8\pm\sqrt{8^2-4\times 3\times 2}}{2\times 3}\]

OpenStudy (anonymous):

\[\frac{8\pm\sqrt{40}}{6}\] and since \(40=4\times 10\) we have \(\sqrt{40}=\sqrt{4}\sqrt{10}=2\sqrt{10}\)

OpenStudy (anonymous):

I have that and understand that step. The step I don't understand is when my teacher went from  and so on